OpenStudy (this):

facotr the qundratic function x^2-20x+100 a,(x+10)^2 b,(x-10)^2 c,(x+10)(x-2) d,(x-10)(x+2)

OpenStudy (anonymous):

x^2-20x+100 (x-10) (x-10)

OpenStudy (anonymous):

worst case you can FOIL all the answers. Only the first two give you 100

OpenStudy (anonymous):

i think b

OpenStudy (anonymous):

\[\checkmark\]

OpenStudy (this):

thanks ya all

OpenStudy (anonymous):

welcome :)

hero (hero):

x^2-20x+100 x^2 - 10x - 10x + 10*10 x(x-10) - 10(x - 10) (x - 10)(x - 10) (x - 10)^2

OpenStudy (anonymous):

x^2-20x+100 think of -10(-10)=100 and-10 + -10 =-20

OpenStudy (anonymous):

so try =(x-10)(x-10) =(x-10)^2